# Ensuring Safety in an Uncertain Environment: Constrained MDPs via Stochastic Thresholds

### v2

@@ G.1 Optimality under Pessimistic Thresholds
Lemma 3 (Policy optimality of pessimistic policies).
Conditioned on the success event $\mathcal{E}$, for any episode $t\in[T^{\prime}]$ and parameter $\gamma\in(0,1]$, under pessimistic thresholds, it holds
$\sum_{t=1}^{T^{\prime}}\tilde{f}_{t}+\lambda_{t}^{T}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{T}y_{\pi^{\star}}\leq\tilde{\mathcal{O}}\left(\left[(1+m\rho)+m\rho\sqrt{SA/\gamma}\right]H^{2}\sqrt{T}\right).$
@@
$\tilde{f}_{t}+\lambda_{t}^{T}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{T}y_{\pi^{\star}}=V_{r}^{\pi^{\star}}(s_{1})-\hat{V}_{\overline{r}}^{\pi_{t}}(s_{1})+\sum_{i=1}^{m}\lambda_{t,i}\left(\overline{\alpha}_{t,i}-\alpha_{i}\right)+\sum_{i=1}^{m}\lambda_{t,i}\left(V_{g_{i}}^{\pi^{\star}}(s_{1})-\hat{V}_{\overline{g}_{i}}^{\pi_{t}}(s_{1})\right).$
@@
$\sum_{t=1}^{T^{\prime}}\tilde{f}_{t}+\lambda_{t}^{T}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{T}y_{\pi^{\star}}$
$=$ $\sum_{t=1}^{T^{\prime}}\left[V_{r+\lambda g}^{\pi^{\star}}(s_{1})-V_{\overline{r}+\lambda\overline{g}}^{\pi_{t}}(s_{1})+\sum_{i=1}^{m}\lambda_{t,i}(\overline{\alpha}_{t,i}-\alpha_{i})\right]$
$=$ $\sum_{t=1}^{T}\sum_{i=1}^{m}\lambda_{t,i}\left[\sum_{h=1}^{H}\overline{\alpha}_{t,i}(s_{h},a_{h})-\alpha_{i}(s_{h},a_{h})\right]$ ($i$)
$+$ $\sum_{t=1}^{T}\sum_{h=1}^{H}\mathbb{E}\Big{[}\langle\hat{Q}_{h}^{t}(s_{h},\cdot),\pi_{h}^{\star}(\cdot\mid s_{h})-\pi_{h}^{t}(\cdot\mid s_{h})\rangle\mid s_{1},\pi^{\star},p\Big{]}$ ($ii$)
$+$ $\sum_{t=1}^{T}\sum_{h=1}^{H}\mathbb{E}\left[-\hat{Q}_{h}^{t}(s_{h},a_{h})+r_{h}(s_{h},a_{h})+\sum_{i=1}^{m}\lambda_{t}g_{i,h}(s_{h},a_{h})+p_{h}(\cdot\mid s_{h},a_{h})\hat{V}_{h+1}^{t}\mid s_{1},\pi^{\star},p\right]$ ($iii$)
@@
$\alpha_{i}(s_{h},a_{h})\leq\overline{\alpha}_{t,i}(s_{h},a_{h}),$
@@
$(i)=\sum_{t=1}^{T}\sum_{i=1}^{m}\lambda_{t,i}\left[\sum_{h=1}^{H}\overline{\alpha}_{t,i}(s_{h},a_{h})-\alpha_{i}(s_{h},a_{h})\right]$
@@
$\sum_{i=1}^{m}\lambda_{t,i}\left[\sum_{h=1}^{H}\overline{\alpha}_{t,i}(s_{h},a_{h})-\alpha_{i}(s_{h},a_{h})\right]$ $\leq\sum_{i=1}^{m}\lambda_{t,i}\left[\sum_{h=1}^{H}\left|\overline{\alpha}_{t,i}(s_{h},a_{h})-\alpha_{i}(s_{h},a_{h})\right|\right]$
$\leq m\max_{i\in[m]}\lambda_{t,i}\left[\sum_{h=1}^{H}\left|\overline{\alpha}_{t,i}(s_{h},a_{h})-\alpha_{i}(s_{h},a_{h})\right|\right]$
@@
By applying Lemmas 4 and 14, and summing over $t$ and $h$, we further have
$\sum_{t=1}^{T}\sum_{i=1}^{m}\lambda_{t,i}\left[\sum_{h=1}^{H}\overline{\alpha}_{t,i}(s_{h},a_{h})-\alpha_{i}(s_{h},a_{h})\right]\leq 2m\rho\sum_{t=1}^{T}\sum_{h=1}^{H}\zeta_{h}^{t}(s_{h},a_{h})$
$\lesssim m\rho\left[\sum_{t=1}^{T}\sum_{h=1}^{H}\mathbb{E}\left[\zeta_{h}^{t}(s_{h},a_{h})\mid s_{1},\pi,p\right]+\sqrt{TH^{3}\text{ln}\frac{1}{\delta}}\right]$
$\lesssim m\rho\left[\sqrt{SAH^{4}T/\gamma}+SAH^{2}+\sqrt{TH^{3}\text{ln}\frac{1}{\delta}}\right]$ (9)
@@
$\sum_{t=1}^{T^{\prime}}\tilde{f}_{t}+\lambda_{t}^{T}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{T}y_{\pi^{\star}}\lesssim\sqrt{H^{4}(1+m\rho)^{2}T}+\sqrt{m^{2}\rho^{2}SAH^{4}T/\gamma}+m\rho SAH^{2}+\sqrt{m^{2}\rho^{2}TH^{3}\text{ln}\frac{1}{\delta}}.$
@@ G.2 Optimality of Optimistic Policies
Lemma 5 (Policy optimality of optimistic policies).
Conditioned on the success event $\mathcal{E}$, for any episode $t\in[T^{\prime}]$, under optimistic thresholds, the following holds:
$\sum_{t=1}^{T^{\prime}}\tilde{f}_{t}+\lambda_{t}^{T}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{T}y_{\pi^{\star}}\leq\tilde{\mathcal{O}}\left((1+m\rho^{\prime})H^{2}\sqrt{T}\right).$
@@
Following the similar initial steps in the proof of Lemma 3, we decompose $\sum_{t=1}^{T^{\prime}}\tilde{f}_{t}+\lambda_{t}^{T}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{T}y_{\pi^{\star}}$ into three terms (I), (II) and (III). The second and third terms (II) and (III) are identical to terms ($ii$) and ($iii$). As previously proved, we have
@@
$(\text{I})=\sum_{t=1}^{T}\sum_{i=1}^{m}\lambda_{t,i}\left[\sum_{h=1}^{H}\underline{\alpha}_{t,i}(s_{h},a_{h})-\alpha_{i}(s_{h},a_{h})\right].$
@@
$\alpha_{i}(s_{h},a_{h})\geq\underline{\alpha}_{t,i}(s_{h},a_{h}),$
@@
$\sum_{t=1}^{T^{\prime}}\tilde{f}_{t}+\lambda_{t}^{T}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{T}y_{\pi^{\star}}\lesssim\tilde{\mathcal{O}}\left((1+m\rho^{\prime})H^{2}\sqrt{T}\right).$
@@ H.1 Proof of Theorem 2
Let $S$, $A$, $m$, $H$, $T$ be the number of states, actions, constraints, horizon length, and episodes, respectively. Define $\rho$ as the Slater gap under pessimistic thresholds, $\gamma$ as the Growing-Window parameter, and $\mathcal{N}$ as the maximum number of non-zero transition probabilities across all state-action pairs. Then with probability at least $1-\delta$, Algorithm 2 under pessimistic thresholds achieves the following regret bounds:
$\mathcal{R}_{T}(r)$ $\leq\tilde{\mathcal{O}}\left(\sqrt{S\mathcal{N}H^{4}T}+\sqrt{(1+m\rho)^{2}SAH^{4}T/\gamma}+(\sqrt{\mathcal{N}}+H)H^{2}SA\right),$
$\mathcal{R}_{T}(g)$ $\leq\tilde{\mathcal{O}}\left(\left(1+\frac{1}{\rho}\right)\left(\sqrt{mS\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H)\sqrt{m}H^{2}SA\right)+m\sqrt{SAH^{4}T/\gamma}\right),$
@@
$\mathcal{L}(\lambda)=$ $\sum_{t=1}^{T^{\prime}}(f_{t}-f_{\text{opt}})+\sum_{t=1}^{T^{\prime}}y_{t}^{T}\lambda$
$\leq$ $\underbrace{\sum_{t=1}^{T^{\prime}}\left(\tilde{f}_{t}+\lambda_{t}^{\top}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{T}y_{\pi^{\star}}\right)}_{(a)}+\underbrace{\sum_{t=1}^{T^{\prime}}(y_{t}-\tilde{y}_{t})^{\top}\lambda}_{(b)}+\underbrace{\sum_{t=1}^{T^{\prime}}(f_{t}-\tilde{f}_{t})}_{(c)}+\underbrace{\frac{1}{2\eta_{\lambda}}\sum_{t=1}^{T^{\prime}}\|\tilde{y}_{t}\|^{2}}_{(d)}+\frac{\eta_{\lambda}}{2}\|\lambda\|_{2}^{2}.$ (14)
@@
Under the pessimistic thresholds, it holds that $\lambda_{t,i}(\overline{\alpha}_{t,i}-\alpha_{i})\geq 0$ for any episode $t$ and constraint $i$. Thus by Lemma 3, we derive
$\sum_{t=1}^{T^{\prime}}\left(\tilde{f}_{t}+\lambda_{t}^{T}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{T}y_{\pi^{\star}}\right)\lesssim\left[(1+m\rho)+m\rho\sqrt{SA/\gamma}\right]H^{2}\sqrt{T}.$
Second term (constraint and threshold estimation error)
In the pessimistic case, it holds that $\alpha_{i}\leq\overline{\alpha}_{t,i}$ for any constraint $i$ and episode $t$. We then obtain
$\left|\sum_{t=1}^{T^{\prime}}(y_{t,i}-\tilde{y}_{t,i})\right|$ $=\left|\sum_{t=1}^{T^{\prime}}\left(\hat{V}_{\overline{g}_{i}}^{\pi_{t}}(s_{1})-V_{g_{i}}^{\pi_{t}}(s_{1})\right)+\sum_{t=1}^{T^{\prime}}\left(\alpha_{i}-\overline{\alpha}_{t,i}\right)\right|$
$\leq\left|\sum_{t=1}^{T^{\prime}}\left(\hat{V}_{\overline{g}_{i}}^{\pi_{t}}(s_{1})-V_{g_{i}}^{\pi_{t}}(s_{1})\right)\right|.$
According to Equation (H.2), we have
$(b)=\sum_{t=1}^{T^{\prime}}(y_{t}-\tilde{y}_{t})^{\top}\lambda\leq$ $\sqrt{\sum_{i=1}^{m}\left(\sum_{t=1}^{T^{\prime}}(y_{t,i}-\tilde{y}_{t,i})\right)^{2}}\|\lambda\|_{2}$
$\leq$ $\sqrt{m}\|\lambda\|_{2}\max_{i\in[m]}\left|\sum_{t=1}^{T^{\prime}}(y_{t,i}-\tilde{y}_{t,i})\right|$
$\leq$ $\sqrt{m}\|\lambda\|_{2}\max_{i\in[m]}\left|\sum_{t=1}^{T^{\prime}}\left(\hat{V}_{\overline{g}_{i}}^{\pi_{t}}(s_{1})-V_{g_{i}}^{\pi_{t}}(s_{1})\right)\right|$
$\leq$ $\tilde{\mathcal{O}}\left(\sqrt{mS\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H)\sqrt{m}H^{2}SA\right)\|\lambda\|_{2}.$
Third term (reward estimation error)
@@
Fourth term (dual update bound)
@@
Substituting these results back into (H.1) and choosing the learning rate $\eta_{\lambda}=\sqrt{\frac{mH^{2}T}{\rho^{2}}}$, we achieve the intermediate bound, for any $0\leq\lambda\leq\rho\mathbf{1}$,
@@
$+\left(\sqrt{mS\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H)\sqrt{m}H^{2}SA\right)\|\lambda\|_{2}$
$+\left[(1+m\rho)+m\rho\sqrt{SA/\gamma}\right]H^{2}\sqrt{T}.$ (15)
@@
$\mathcal{R}_{T}(r)\leq\tilde{\mathcal{O}}\left(\sqrt{S\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H)H^{2}SA+\left[(1+m\rho)+m\rho\sqrt{SA/\gamma}\right]H^{2}\sqrt{T}\right).$
@@
For constraint $i\in[m]$, set the vector $\lambda_{i}=\rho e_{i}$ if $[\sum_{t=1}^{T^{\prime}}y_{t,i}]_{+}>0$ and 00 otherwise in (H.1), where $e_{i}$ denotes the standard basis vector and the Slater gap $\rho$ is defined in Assumption 2. This also guarantees $\|\lambda_{i}\|_{2}^{2}\leq\rho^{2}$. We then have
@@
$\lesssim$ $\left(1+\rho\right)\left(\sqrt{mS\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H)\sqrt{m}H^{2}SA\right)+\left[(1+m\rho)+m\rho\sqrt{SA/\gamma}\right]H^{2}\sqrt{T}:=\theta(T).$ (16)
Due to the linearity of both the reward and constraints, combined with the convexity of the occupancy-measure set, we leverage Lemmas 23 and 24 and then derive the following result for any $T^{\prime}\in[T]$:
@@
$\mathcal{R}_{T}(g)\leq\tilde{\mathcal{O}}\left(\left(1+\frac{1}{\rho}\right)\left(\sqrt{mS\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H)\sqrt{m}H^{2}SA\right)+m\sqrt{SAH^{4}T/\gamma}\right).$

### v3

@@ G.1 Optimality under Pessimistic Thresholds
Lemma 3 (Policy optimality under pessimistic thresholds).
Conditioned on the success event $\mathcal{E}$, for any episode $t\in[T^{\prime}]$ and parameter $\gamma\in(0,1]$, if we use pessimistic thresholds, then
$\sum_{t=1}^{T^{\prime}}\tilde{f}_{t}+\lambda_{t}^{\top}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{\top}y_{\pi^{\star}}\leq\tilde{\mathcal{O}}\left(\sqrt{H^{4}(1+m\rho)^{2}T}+\sqrt{m^{2}\rho^{2}SAH^{2}T/\gamma}\right).$
@@
$\tilde{f}_{t}+\lambda_{t}^{\top}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{\top}y_{\pi^{\star}}=V_{r}^{\pi^{\star}}(s_{1})-\hat{V}_{\overline{r}}^{\pi_{t}}(s_{1})+\sum_{i=1}^{m}\lambda_{t,i}\left(\overline{\alpha}_{i}^{t}-\alpha_{i}\right)+\sum_{i=1}^{m}\lambda_{t,i}\left(V_{g_{i}}^{\pi^{\star}}(s_{1})-\hat{V}_{\overline{g}_{i}}^{\pi_{t}}(s_{1})\right).$
@@
$\sum_{t=1}^{T^{\prime}}\tilde{f}_{t}+\lambda_{t}^{\top}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{\top}y_{\pi^{\star}}$
$=$ $\sum_{t=1}^{T^{\prime}}\left[V_{r+\lambda g}^{\pi^{\star}}(s_{1})-V_{\overline{r}+\lambda\overline{g}}^{\pi_{t}}(s_{1})+\sum_{i=1}^{m}\lambda_{t,i}(\overline{\alpha}_{i}^{t}-\alpha_{i})\right]$
$=$ $\sum_{t=1}^{T}\sum_{i=1}^{m}\lambda_{t,i}\left[\sum_{h=1}^{H}\overline{\alpha}_{i}^{t}(s_{h},a_{h})-\alpha_{i}(s_{h},a_{h})\right]$ (ii)
$+$ $\sum_{t=1}^{T}\sum_{h=1}^{H}\mathbb{E}\Big[\langle\hat{Q}_{h}^{t}(s_{h},\cdot),\pi_{h}^{\star}(\cdot\mid s_{h})-\pi_{h}^{t}(\cdot\mid s_{h})\rangle\mid s_{1},\pi^{\star},p\Big]$ (iiii)
$+$ $\sum_{t=1}^{T}\sum_{h=1}^{H}\mathbb{E}\left[-\hat{Q}_{h}^{t}(s_{h},a_{h})+r_{h}(s_{h},a_{h})+\sum_{i=1}^{m}\lambda_{t}g_{i,h}(s_{h},a_{h})+p_{h}(\cdot\mid s_{h},a_{h})\hat{V}_{h+1}^{t}\mid s_{1},\pi^{\star},p\right]$ (iiiiii)
@@
$\alpha_{i}(s_{h},a_{h})\leq\overline{\alpha}_{i}^{t}(s_{h},a_{h}),$
@@
$(i)=\sum_{t=1}^{T}\sum_{i=1}^{m}\lambda_{t,i}\left[\sum_{h=1}^{H}\overline{\alpha}_{i}^{t}(s_{h},a_{h})-\alpha_{i}(s_{h},a_{h})\right]$
@@
$\sum_{i=1}^{m}\lambda_{t,i}\left[\sum_{h=1}^{H}\overline{\alpha}_{i}^{t}(s_{h},a_{h})-\alpha_{i}(s_{h},a_{h})\right]$ $\leq\sum_{i=1}^{m}\lambda_{t,i}\left[\sum_{h=1}^{H}\left|\overline{\alpha}_{i}^{t}(s_{h},a_{h})-\alpha_{i}(s_{h},a_{h})\right|\right]$
$\leq m\max_{i\in[m]}\lambda_{t,i}\left[\sum_{h=1}^{H}\left|\overline{\alpha}_{i}^{t}(s_{h},a_{h})-\alpha_{i}(s_{h},a_{h})\right|\right]$
@@
By applying Lemmas 12 and 13, and summing over $t$ and $h$, we further have
$\sum_{t=1}^{T}\sum_{i=1}^{m}\lambda_{t,i}\left[\sum_{h=1}^{H}\overline{\alpha}_{i}^{t}(s_{h},a_{h})-\alpha_{i}(s_{h},a_{h})\right]\leq 2m\rho\sum_{t=1}^{T}\sum_{h=1}^{H}\zeta_{h}^{t}(s_{h},a_{h})$
$\lesssim m\rho\left[\sum_{t=1}^{T}\sum_{h=1}^{H}\mathbb{E}\left[\zeta_{h}^{t}(s_{h},a_{h})\mid s_{1},\pi,p)\right]+\sqrt{TH^{3}\text{ln}\frac{1}{\delta}}\right]$
$\lesssim m\rho\left[\sqrt{SAH^{2}T/\gamma}+SAH+\sqrt{TH\text{ln}\frac{1}{\delta}}\right]$ (8)
@@
$\sum_{t=1}^{T^{\prime}}\tilde{f}_{t}+\lambda_{t}^{\top}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{\top}y_{\pi^{\star}}\lesssim\sqrt{H^{4}(1+m\rho)^{2}T}+\sqrt{m^{2}\rho^{2}SAH^{2}T/\gamma}+m\rho SAH+\sqrt{m^{2}\rho^{2}TH\text{ln}\frac{1}{\delta}}.$
@@ G.2 Optimality of Optimistic Policies
Lemma 5 (Policy optimality with optimistic thresholds).
Conditioned on the success event $\mathcal{E}$, for any episode $t\in[T^{\prime}]$, if we adopt the optimistic thresholds, then
$\sum_{t=1}^{T^{\prime}}\tilde{f}_{t}+\lambda_{t}^{\top}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{\top}y_{\pi^{\star}}\leq\tilde{\mathcal{O}}\left((1+m\rho^{\prime})H^{2}\sqrt{T}\right).$
@@
Following the similar initial steps in the proof of Lemma 3, we decompose $\sum_{t=1}^{T^{\prime}}\tilde{f}_{t}+\lambda_{t}^{\top}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{\top}y_{\pi^{\star}}$ into three terms (I), (II) and (III). The second and third terms (II) and (III) are identical to terms ($ii$) and (iiiiii). As previously proved, we have
@@
$(\text{I})=\sum_{t=1}^{T}\sum_{i=1}^{m}\lambda_{t,i}\left[\sum_{h=1}^{H}\underline{\alpha}_{i}^{t}(s_{h},a_{h})-\alpha_{i}(s_{h},a_{h})\right].$
@@
$\alpha_{i}(s_{h},a_{h})\geq\underline{\alpha}_{i}^{t}(s_{h},a_{h}),$
@@
$\sum_{t=1}^{T^{\prime}}\tilde{f}_{t}+\lambda_{t}^{\top}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{\top}y_{\pi^{\star}}\lesssim\tilde{\mathcal{O}}\left((1+m\rho^{\prime})H^{2}\sqrt{T}\right).$
@@ H.1 Proof of Theorem 2
Let $S$ be the number of states, $A$ be the number of actions, $H$ be the horizon length, $m$ be the number of constraints, $\rho$ be the Slater gap under pessimistic thresholds, $\gamma$ be the Growing-Window parameter, $T$ be the number of episodes and $\mathcal{N}$ denote the maximum number of non-zero transition probabilities across all state-action pairs. Then with probability $1-\delta$, Algorithm 2 under pessimistic thresholds obtains the following regret bounds:
$\mathcal{R}_{T}(r)$ $\leq\tilde{\mathcal{O}}\left(\sqrt{S\mathcal{N}H^{4}T}+\sqrt{H^{4}(1+m\rho)^{2}T}+\sqrt{m^{2}\rho^{2}SAH^{2}T/\gamma}+(\sqrt{\mathcal{N}}+H)H^{2}SA\right),$
$\mathcal{R}_{T}(g)$ $\leq\tilde{\mathcal{O}}\left(\frac{1}{\rho}\left(\sqrt{S\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H)H^{2}SA\right)+m\sqrt{SAH^{4}T/\gamma}\right),$
@@
$\mathcal{L}(\lambda)=$ $\sum_{t=1}^{T^{\prime}}(f_{t}-f_{\text{opt}})+\sum_{t=1}^{T^{\prime}}y_{t}^{\top}\lambda$
$\leq$ $\sum_{t=1}^{T^{\prime}}(y_{t}-\tilde{y}_{t})^{\top}\lambda+\sum_{t=1}^{T^{\prime}}\left(\tilde{f}_{t}+\lambda_{t}^{\top}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{\top}y_{\pi^{\star}}\right)+\sum_{t=1}^{T^{\prime}}(f_{t}-\tilde{f}_{t})+\frac{1}{2\eta_{\lambda}}\sum_{t=1}^{T^{\prime}}\|\tilde{y}_{t}\|^{2}+\frac{\eta_{\lambda}}{2}\|\lambda\|_{2}^{2}.$
Under the pessimistic thresholds, we have $y_{t}\leq\tilde{y}_{t}$ at episode $t$, thus
$\sum_{t=1}^{T^{\prime}}(y_{t}-\tilde{y}_{t})\lambda\leq 0,$
leading us to simplify the above inequality as
$\mathcal{L}(\lambda)\leq\underbrace{\sum_{t=1}^{T^{\prime}}\left(\tilde{f}_{t}+\lambda_{t}^{\top}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{\top}y_{\pi^{\star}}\right)}_{(a)}+\underbrace{\sum_{t=1}^{T^{\prime}}(f_{t}-\tilde{f}_{t})}_{(b)}+\underbrace{\frac{1}{2\eta_{\lambda}}\sum_{t=1}^{T^{\prime}}\|\tilde{y}_{t}\|^{2}}_{(c)}+\frac{\eta_{\lambda}}{2}\|\lambda\|_{2}^{2}.$ (13)
@@
Under the pessimistic thresholds, it holds that $\lambda_{t,i}(\overline{\alpha}_{i}^{t}-\alpha_{i})\geq 0$ for any episode $t$ and constraint $i$. Thus by Lemma 3, we derive
$\sum_{t=1}^{T^{\prime}}\left(\tilde{f}_{t}+\lambda_{t}^{\top}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{\top}y_{\pi^{\star}}\right)\lesssim\sqrt{H^{4}(1+m\rho)^{2}T}+\sqrt{m^{2}\rho^{2}SAH^{2}T/\gamma}.$
Second term (reward estimation error)
@@
Third term (dual update bound)
@@
Substituting these results back into (13) and choosing the learning rate $\eta_{\lambda}=\sqrt{\frac{mH^{2}T}{\rho^{2}}}$, we achieve the intermediate bound, for any $0\leq\lambda\leq\rho\mathbf{1}$,
@@
$+\sqrt{H^{4}(1+m\rho)^{2}T}+\sqrt{m^{2}\rho^{2}SAH^{2}T/\gamma}.$ (14)
@@
$\mathcal{R}_{T}(r)\leq\tilde{\mathcal{O}}\left(\sqrt{S\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H)H^{2}SA+\sqrt{H^{4}(1+m\rho)^{2}T}+\sqrt{m^{2}\rho^{2}SAH^{2}T/\gamma}\right).$
@@
For constraint $i\in[m]$, set the vector $\lambda_{i}=\rho e_{i}$ if $\sum_{t=1}^{T^{\prime}}y_{t,i}]_{+}>0$ and 0 otherwise in (H.1), where $e_{i}$ denotes the standard basis vector and the Slater gap $\rho$ is defined in Assumption 1. This also guarantees $\|\lambda_{i}\|_{2}^{2}\leq\rho^{2}$. We then have
@@
$\lesssim$ $\left(\sqrt{S\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H)H^{2}SA\right)+\sqrt{H^{4}(1+m\rho)^{2}T}+\sqrt{m^{2}\rho^{2}SAH^{2}T/\gamma}:=\theta(T).$ (15)
Due to the linearity of both the reward and constraints, combined with the convexity of the occupancy-measure set, we leverage Lemmas 22 and 23 and then derive the following result for any $T^{\prime}\in[T]$:
@@
$\mathcal{R}_{T}(g)\leq\tilde{\mathcal{O}}\left(\frac{1}{\rho}\left(\sqrt{S\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H)H^{2}SA\right)+m\sqrt{SAH^{4}T/\gamma}\right).$